OpenStudy (anonymous):

What is the perimeter of a rectangle with a length of 7 3/4 inches and a width of 4 1/3 inches?

OpenStudy (igreen):

Formula: \(P = 2l + 2w\)

OpenStudy (igreen):

Plug in what we know: \(P = 2(7 \dfrac{3}{4}) + 2(4 \dfrac{1}{3})\) Simplify: \(P = 15 \dfrac{1}{2} + 8 \dfrac{2}{3}\) Can you add that? @Micahc

OpenStudy (anonymous):

ok would it be 24 1/6 ? @iGreen

OpenStudy (igreen):

Yep, you got it!